OpenStudy (anonymous):

M is the midpoint of segment AB, AM=2x+4 and AB=12x+4. What is the value of MB

OpenStudy (anonymous):

If M is the midpoint, then AM = MB. Also, AM + MB = AB. By substution, therefore, AM + AM = AB: (2x + 4) + (2x + 4) = 12x + 4 4x + 8 = 12x + 4 4 = 8x 1/2 = x AM = 2(1/2) + 4 = 5 MB = AM = 5